OpenStudy (anonymous):

Find the area of an equilateral triangle (regular 3-gon) with the given measurement. 6-inch apothem

OpenStudy (anonymous):

I've done this problem before, does it show the following... "Find the area of an equilateral triangle (regular 3-gon) with the given measurement. 6-inch apothem A = sq. in."?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

6-inch apothem, apothem a = 6 an equilateral triangle has 6 30-60-90 triangles each with the shortest leg being the apothem, length = a = 6 find hypotenuse of each of the 6 triangles: angle opposite to a is 30 degrees, sin 30 = opposite/hypotenuse = a/hyp = 6/hyp 0.5 = 6/hyp 0.5hyp = 6 hyp = 12 inches Can you tell me, how to find adjacent of each of the 6 triangles?